# Post-hoc Popularity Bias Correction in GNN-based Collaborative Filtering

### v1

@@ 5. Experiments
In this section, we evaluate the performance of our PPD method by addressing the following research questions:
@@
Table 2. A comparison of different methods across datasets and evaluation metrics using LightGCN as the backbone. The best-performing method is shown in bold, and the second-best method is underlined. Performance improvement (%) is given compared to the best baseline.
@@
Recall NDCG HR Recall NDCG HR Recall NDCG HR
Post-hoc PPD 0.0988 0.2899 0.9043 0.2407 0.1957 0.5871 0.1505 0.0695 0.2237
DAP (Chen et al., 2024) 0.0289 0.0284 0.3007 0.1607 0.0777 0.4555 0.1416 0.0647 0.2150
None LightGCN (He et al., 2020) 0.0055 0.0043 0.0709 0.1572 0.0773 0.4626 0.1403 0.0640 0.2136
In-training APDA (Zhou et al., 2023) 0.0097 0.0077 0.1057 0.1435 0.0727 0.4413 0.1392 0.0627 0.2081
NAVIP (Kim et al., 2022) 0.0006 0.0039 0.0759 0.1413 0.0730 0.4377 0.1439 0.0648 0.2161
MACR (Wei et al., 2021) 0.0104 0.0107 0.1872 0.1002 0.0451 0.3345 0.0089 0.0033 0.0181
IPSCN (Joachims et al., 2017) 0.0010 0.0060 0.0574 0.2296 0.1933 0.5765 0.1309 0.0589 0.1977
SAM-REG (Boratto et al., 2021) 0.0018 0.0133 0.1121 0.2206 0.1892 0.5730 0.1226 0.0551 0.1847
PPAC (Ning et al., 2024) 0.0345 0.0483 0.6177 0.1444 0.0716 0.3915 0.1462 0.0657 0.2178
Improvement (%) +186.4% +500.2% +46.4% +4.8% +1.2% +1.8% +2.9% +5.8% +2.7%
5.1. Experimental Setup
Datasets. We use three real-world datasets: KuaiRec (Gao et al., 2022a), Coat (Schnabel et al., 2016), and Yahoo! R3 (Marlin and Zemel, 2009). We choose these datasets because they provide unbiased test data, which allows us unbiased evaluation of methods. KuaiRec, from Kuaishou short-video logs, provides a fully observed user–item matrix for unbiased evaluation, with nearly all users rating every video, while the training set remains sparse. The test set includes ratings from $1,411$ users on $3,327$ items. Following (Gao et al., 2022a), an interaction is considered positive if a user watches a video for more than twice its duration. Coat contains customer ratings from an online coat store, where each user rated $24$ self-selected items in the training set, and the unbiased test set includes ratings from $16$ randomly shown items per user. Yahoo! R3 dataset contains over $300,000$ self-selected user–song ratings by $15,400$ users in the training set. Additionally, a separate unbiased test set is collected by asking $5,400$ users to provide ratings for $10$ randomly presented songs. Ratings are given on a $5$-point scale, and for our experiments, we consider ratings of $3$ or higher as positive feedback. Table 1 presents a detailed overview of the datasets.
@@ 5.1. Experimental Setup
Evaluation metrics. Following the all-ranking protocol (Krichene and Rendle, 2020), we rank all items per user, excluding those with positive feedback in the training set. Recommendation performance is measured using Recall@$k$, NDCG@$k$, and Hit Ratio (HR@$k$) with $k=20$ by default. $k$ denotes the position in a ranked list up to which we evaluate a model’s performance. However, the KuaiRec dataset presents a special case. Its test set is constructed from a dense subset of the user-item matrix (Gao et al., 2022a), rather than being a random sample from the entire space. As a result, the traditional all-ranking method is not applicable. Instead, for KuaiRec dataset, we evaluate by ranking only the $3,327$ items that are fully exposed.
@@
Baselines. We evaluate the performance of our method relative to popularity bias correction approaches in CF, including post-hoc method DAP (Chen et al., 2024), interaction-weighting techniques applied during aggregation like APDA (Zhou et al., 2023) and NAVIP (Kim et al., 2022), causal inference-based approaches such as MACR (Wei et al., 2021) and PPAC (Ning et al., 2024), propensity weighting methods like IPSCN (Gruson et al., 2019), and a regularization-based approach SAM-REG (Boratto et al., 2021). We also include LightGCN (He et al., 2020) as a standard GNN-based CF backbone without any debiasing, using it as the backbone for all baselines as well as for our method. These baselines cover the major classes of debiasing strategies, while detailed descriptions of each method are provided in the Appendix B.1.
Implementation details. Since LightGCN (He et al., 2020) is the most widely used backbone for GNN-based CF, we adopt it for all baselines and our method, unless otherwise specified. The model is implemented in PyTorch and optimized using Adam. We apply early stopping with a patience of $100$ epochs, stopping if Recall@$20$ on the validation set does not improve for $100$ consecutive epochs. This setting allows the model to achieve its best performance while maintaining the early stopping criterion. We tune the learning rate among $1{e}^{-3},1{e}^{-4},1{e}^{-5}$ and the regularization coefficient among $0.0,1{e}^{-3},1{e}^{-5}$. Batch size is 2,0482{,}048 and embedding dimension is $256$. All other parameters follow the default settings of LightGCN. Unless stated otherwise, we use a two-layer GCN that integrates up to two-hop neighbors. Popularity scores ($b_{ui}$) are computed by min–max normalizing global preference ($p_{i}$) and personalized preference ($r_{ui}$). Hyperparameters for our method are tuned with popularity penalty coefficient $\beta\in\{0.0,0.1,0.2,0.3\}$ and preference centroid coefficient $\phi\in\{0.0,0.25,0.50,0.75,1.0\}$. $\beta$ is restricted up to 0.3 to avoid over-penalizing popularity, which could suppress preference signals and weaken the estimation of personalized preference. The small constant $\epsilon$ in equations (8) and (9) is set to $1{e}^{-8}$. For unbiased validation, we split the unbiased test data of each original dataset, using one-third for validation and the remaining two-thirds for testing. We use the dot product as the ${sim}(\cdot)$ function in equations (5) and (6). To validate our method on other GNN-based CF models, we also use SGL (Wu et al., 2021) as a backbone (see RQ4). SGL augments LightGCN with a self-supervised contrastive loss. Following (Zhang et al., 2023b), in our experiments, we set the contrastive loss temperature to $0.15$ and its weight to $0.2$, while keeping all other settings at their default LightGCN values.
@@ 5.2. Results
Performance comparison between our method and the baselines (RQ1). In response to RQ1, we evaluate whether our method outperforms existing popularity bias correction approaches. Table 2 compares PPD with baselines, where in-training methods correct bias during training and post-hoc methods do so after training. PPD consistently achieves the best results across all metrics and datasets. On KuaiRec, where popularity bias can be severe (Figure 1), PPD outperforms the strongest baseline (PPAC) with relative improvements of 186.4% in Recall@20, 500.2% in NDCG@20, and 46.4% in HR@20, highlighting its effectiveness in highly skewed data. On Coat, with milder popularity (Figure 1), PPD outperforms the strongest baseline IPSCN by 4.8% in Recall@20, 1.2% in NDCG@20, and 1.8% in HR@20, remaining competitive in less biased scenarios. On sparse Yahoo! R3 (Table 1), PPD exceeds all baselines, improving 2.9%, 5.8%, and 2.7% over PPAC. PPD also outperforms the post-hoc baseline DAP (Chen et al., 2024), as it estimates interaction-level popularity for debiasing by considering both global and personalized preferences, rather than clustering similar nodes and relying on simple node degree information to remove bias from embeddings.
These results demonstrate several key insights. First, the substantial improvements on KuaiRec emphasize that PPD is particularly effective in settings where popularity can be most prominent, confirming its robustness to extreme skew in interaction frequency. Second, the consistent improvements on Coat indicate that PPD also generalizes well to datasets with less pronounced bias. Finally, even on Yahoo! R3, where extremely sparse interactions can make debiasing particularly difficult due to fewer preference signals in the data, PPD yields consistent improvements, suggesting its ability to correct for bias in sparser datasets. Together, these findings show that PPD consistently outperforms the baselines across datasets with varying levels of skewness and sparsity.
Popular and niche item performance improvement (RQ2). RQ2 evaluates performance on popular (head) and niche (tail) items, examining whether improvements on one group come at the expense of the other. Popularity bias correction can create such a trade-off, where enhancing tail-item performance may degrade head-item accuracy (Zhou et al., 2023). To investigate this, we divide items into the top 20% most frequently interacted ones (popular) and the remaining 80% (niche) based on the training data, and evaluate performance on these subsets across all three datasets. This allows assessment of whether PPD balances recommendation quality for both head and tail items. As shown in Table 3, PPD consistently achieves the best performance across both groups. On KuaiRec, PPD improves head performance by +186.7% in Recall@20 and +492.3% in NDCG@20 compared to the best baseline. For niche items, where all methods struggle due to very skewed data in tails, PPD still shows the best performance, achieving relative improvements of +233.3% in Recall@20 and +400% in NDCG@20 compared to best baseline. The overall low scores on KuaiRec tails stem from the extremely imbalanced distribution, where most niche items may have only a few interactions, making them inherently difficult to recommend. On Coat, PPD also improves both head and niche groups simultaneously. For head items, it yields +2.7% Recall@20 and +0.94% NDCG@20 over the best baseline. For niche items, it further improves by +1.24% Recall@20 and +1.09% NDCG@20, showing that gains on tail recommendations do not compromise head performance. On Yahoo! R3, PPD improves head recommendations with +7.1% Recall@20 and +12.0% NDCG@20 compared to best baseline, while also improving tail performance with +2.33% Recall@20, though NDCG shows a slight drop (–1.5%) compared to best baseline. Across all datasets, PPD achieves the highest HR@20 compared to the baselines. We show the results for HR@20 in Table 6 in the Appendix.
In summary, our method enhances recommendations for tail items, demonstrating its ability to promote less popular or niche items while maintaining strong accuracy for head items. PPD explicitly removes popularity-driven components from embeddings while retaining preference-driven signals through interaction-level popularity scores. This ensures that head items are recommended when they genuinely reflect user interests, and tail items are recommended when aligned with user preferences. As a result, the typical head–tail trade-off is avoided: head accuracy is preserved, and tail performance is improved.
@@
Method KuaiRec Coat Yahoo! R3
Recall NDCG HR Recall NDCG HR Recall NDCG HR
PPD 0.1352 0.3612 0.9978 0.2437 0.2176 0.6121 0.1381 0.0601 0.2031
Improvement (%) +1017.4% +1811.1% +287.6% +5.8% +7.3% +5.6% +8.1% +3.1% +6.4%
Method KuaiRec Coat Yahoo! R3
Recall NDCG HR Recall NDCG HR Recall NDCG HR
PPD 0.0988 0.2899 0.9043 0.2407 0.1957 0.5871 0.1505 0.0695 0.2237
@@
Effect of GNN layer depth on model performance (RQ3). We investigate how varying the number of GNN layers influences the performance of the PPD method. Figure 2 demonstrates that the effect of increasing the number of layers. On the KuaiRec dataset, where popularity bias is most severe, LightGCN reaches peak performance with two layers but suffers degradation as more layers are added, indicating that deeper propagation amplifies popularity bias and over-smoothing. In contrast, PPD continues to benefit from additional layers, suggesting that its debiasing mechanism can mitigate bias amplification while leveraging multi-hop information. On the Coat dataset, where popularity bias is less pronounced, both models achieve their best performance with two layers, and deeper propagation beyond this point may introduce noise rather than useful signals, reducing accuracy. For Yahoo! R3, PPD achieves optimal results with two layers, while LightGCN slightly peaks at three layers, reflecting that in moderately biased and sparse settings, LightGCN can still gain from slightly deeper propagation. Importantly, across all datasets and all depths, PPD consistently outperforms LightGCN, demonstrating robustness and resilience to over-smoothing and bias mitigation. These findings highlight that the optimal number of layers depends on the dataset, and PPD maintains stronger performance by mitigating popularity bias across different depths. We include the results for Recall@20 here, and in Figure 5 in the Appendix, we show the corresponding NDCG@20 results for the three datasets.
Performance of PPD with SGL as backbone (RQ4). To validate the generalizability of our method, we apply PPD to another widely used GNN-based CF backbone, Self-supervised Graph Learning (SGL) (Wu et al., 2021). SGL augments LightGCN with self-supervised contrastive learning, generating multiple graph views via node dropout, edge dropout, and random walks. By maximizing agreement between views of the same node, SGL enhances representation robustness and recommendation accuracy. Evaluating PPD on SGL tests whether our debiasing strategy benefits other GNN-based CF backbones. Table 4 shows results on KuaiRec, Coat, and Yahoo! R3. On KuaiRec, PPD achieves $1017.4\%$ Recall@20, $1811.1\%$ NDCG@20, and $287.6\%$ HR@20 over SGL. For Coat, improvements are $5.8\%$ Recall@20, $7.3\%$ NDCG@20, and $5.6\%$ HR@20, while on Yahoo! R3, gains are $8.1\%$ Recall@20, $3.1\%$ NDCG@20, and $6.4\%$ HR@20. These results highlight two key findings. First, PPD shows substantial improvements on datasets with strong popularity bias, such as KuaiRec, where SGL alone struggles to capture unbiased signals. Second, even on less biased datasets like Coat and Yahoo! R3, PPD still yields consistent improvements across all evaluation metrics. This demonstrates that PPD is not limited to a specific backbone but can be effectively applied on top of stronger backbones like SGL, further boosting recommendation accuracy through mitigating popularity bias.
(a) KuaiRec dataset
(b) Coat dataset
(c) Yahoo! R3 dataset
(a) KuaiRec dataset
(b) Coat dataset
(c) Yahoo! R3 dataset
Impact of incorporating personalized preference into popularity score estimation (RQ5). In RQ5, we examine the role of personalized preference ($r_{ui}$) in estimating the popularity score ($b_{ui}$). To this end, we introduce a variant, PPD w/o-PP, which computes the popularity score solely based on global preference ($p_{i}$) without accounting for user-specific preferences with items. The results in Table 5 show that PPD consistently outperforms PPD w/o-PP across all datasets and evaluation metrics. This comparison reveals that relying only on global preference can be insufficient, as popular items may still align with true user interests. Without personalized preference, the popularity score ($b_{ui}$) risks suppressing relevant popular items, conflating popularity-driven interactions with genuine user preferences. By removing personalized preference from global preference, the popularity score captures only the residual effect of popularity. These results demonstrate that accounting for personalized preference in estimating the popularity score is important, as it enables PPD to more accurately estimate the true popularity effect, enabling better performance across datasets.
Effect of the popularity penalty coefficient ($\bm{\beta}$) on PPD (RQ7). In RQ7, we evaluate the effect of the popularity penalty coefficient ($\beta$), which controls the strength of penalizing global preference when estimating personalized preferences. As shown in Figure 4, performance varies across datasets and values of $\beta$. On KuaiRec, Recall@20 improves as $\beta$ increases up to 0.2 but declines when raised to 0.3, indicating that moderate global preference penalization yields the best personalized preference and, consequently, the highest accuracy. On Coat, the highest performance occurs at $\beta=0.1$, with larger values leading to reduced Recall@20. In contrast, Yahoo! R3 achieves its peak performance at $\beta=0.3$. These results highlight two key insights. First, without any penalty ($\beta=0.0$), the model cannot achieve its best performance, demonstrating the necessity of incorporating a popularity penalty in estimating personalized preferences. Second, the optimal value of $\beta$ is dataset-dependent, indicating that it should be tuned for each dataset to achieve optimal performance.

### v2

@@ 5. Experiments
In this section, we evaluate the performance of our PPD method by formulating the following research questions:
@@
5.1. Experimental setup
Datasets. We use three real-world datasets: KuaiRec (Gao et al., 2022b), Coat (Schnabel et al., 2016), and Yahoo! R3 (Marlin and Zemel, 2009). We choose these datasets because they provide unbiased test data, which allows us unbiased evaluation of methods. KuaiRec, from Kuaishou short-video logs, provides a fully observed user–item matrix for unbiased evaluation, with nearly all users rating every video, while the training set remains sparse. The test set includes ratings from $1,411$ users on $3,327$ items. Following (Gao et al., 2022b), an interaction is considered positive if a user watches a video for more than twice its duration. Coat contains customer ratings from an online coat store, where each user rated $24$ self-selected items in the training set, and the unbiased test set includes ratings from $16$ randomly shown items per user. Yahoo! R3 dataset contains over $300,000$ self-selected user–song ratings by $15,400$ users in the training set. Additionally, a separate unbiased test set is collected by asking $5,400$ users to provide ratings for $10$ randomly presented songs. Ratings are given on a $5$-point scale, and for our experiments, we consider ratings of $3$ or higher as positive feedback. Table 1 presents a detailed overview of the datasets. Although evaluation on other datasets is possible, they generally do not provide unbiased test data. Those datasets construct test sets by sampling from user interactions in the training data (Wei et al., 2021), which still follow a popularity-dominated distribution and prevent unbiased evaluation. Therefore, we use the three mentioned datasets, which provide unbiased test data under random or full-exposure settings, varying sizes, sparsity, and popularity, covering diverse real-world scenarios.
@@ 5.1. Experimental setup
Table 2. A comparison of different methods across datasets and evaluation metrics using LightGCN as the backbone. The best-performing method is shown in bold, and the second-best method is underlined. Performance improvement (%) is given compared to the best baseline for all evaluation metrics. We additionally report the runtime (in minutes) for each method and the corresponding runtime improvement of PPD in terms of time speedup compared to the best-performing baseline.
@@
Recall NDCG HR Runtime Recall NDCG HR Runtime Recall NDCG HR Runtime
Post-hoc PPD 0.0988 0.2899 0.9043 112.31 0.2407 0.1957 0.5871 0.11 0.1505 0.0695 0.2237 3.35
DAP (Chen et al., 2024) 0.0289 0.0284 0.3007 260.13 0.1607 0.0777 0.4555 0.28 0.1416 0.0647 0.2150 32.25
None LightGCN (He et al., 2020) 0.0055 0.0043 0.0709 115.42 0.1572 0.0773 0.4626 2.97 0.1403 0.0640 0.2136 13.83
In-training APDA (Zhou et al., 2023) 0.0097 0.0077 0.1057 832.52 0.1435 0.0727 0.4413 0.68 0.1392 0.0627 0.2081 31.58
NAVIP (Kim et al., 2022) 0.0006 0.0039 0.0759 121.34 0.1413 0.0730 0.4377 0.35 0.1439 0.0648 0.2161 13.30
MACR (Wei et al., 2021) 0.0104 0.0107 0.1872 119.23 0.1002 0.0451 0.3345 1.17 0.0089 0.0033 0.0181 4.03
IPSCN (Joachims et al., 2017) 0.0010 0.0060 0.0574 118.32 0.2296 0.1933 0.5765 3.15 0.1309 0.0589 0.1977 5.12
SAM-REG (Boratto et al., 2021) 0.0018 0.0133 0.1121 132.12 0.2206 0.1892 0.5730 3.75 0.1226 0.0551 0.1847 7.14
PPAC (Ning et al., 2024) 0.0345 0.0483 0.6177 260.21 0.1444 0.0716 0.3915 0.26 0.1462 0.0657 0.2178 11.52
Improvement +186.4% +500.2% +46.4% 2.32× +4.8% +1.2% +1.8% 28.64× +2.9% +5.8% +2.7 3.44×
@@
Evaluation metrics. Following the all-ranking protocol (Krichene and Rendle, 2020), we rank all items per user, excluding those with positive feedback in the training set. Recommendation performance is measured using Recall@$k$, NDCG@$k$, and Hit Ratio (HR@$k$) with $k=20$ by default, where $k$ is the evaluation cutoff. The KuaiRec dataset presents a special case. Its test set is constructed from a dense subset of the user-item matrix (Gao et al., 2022b), rather than being a random sample from the entire space. As a result, the traditional all-ranking protocol is not applicable. Instead, for KuaiRec dataset, we evaluate by ranking only the $3,327$ items that are fully exposed.
@@
Baselines. We evaluate the performance of PPD relative to popularity bias correction methods in CF, including post-hoc method DAP (Chen et al., 2024), interaction-weighting methods applied during aggregation like APDA (Zhou et al., 2023) and NAVIP (Kim et al., 2022), causal inference-based approaches such as MACR (Wei et al., 2021) and PPAC (Ning et al., 2024), IPW methods like IPSCN (Gruson et al., 2019), and a regularization-based approach SAM-REG (Boratto et al., 2021). We include LightGCN (He et al., 2020) as a standard GNN-based CF backbone without any debiasing, using it as the backbone for all baselines and for our method. These baselines cover all the major classes of debiasing strategies, while detailed descriptions of baselines are provided in the Appendix B.1. Since PPD focuses on pure CF using only interaction data, without any user or item features or side information, we include baselines that are adaptable to this setting. We provide the detailed implementation details in the Appendix B.2. Our GitHub code is publicly available in the provided link [^1] Code is available at https://github.com/edgeslab/Posthoc_Popularity_Debiasing.
@@ 5.2. Results
Performance comparison between our method and the baselines (RQ1). In response to RQ1, we evaluate whether PPD outperforms existing popularity bias correction approaches. Table 2 compares PPD with baselines, where in-training methods correct bias during training and post-hoc methods do so after training. PPD consistently achieves the best results across all metrics. On KuaiRec, where popularity bias can be severe (Figure 1), PPD outperforms the strongest baseline (PPAC) with relative improvements of 186.4% in Recall@20, 500.2% in NDCG@20, and 46.4% in HR@20, highlighting its effectiveness in highly skewed data. On Coat, with milder popularity (Figure 1), PPD outperforms the strongest baseline IPSCN by 4.8% in Recall@20, 1.2% in NDCG@20, and 1.8% in HR@20, remaining competitive in less biased scenarios. On sparse Yahoo! R3 (Table 1), PPD exceeds all baselines, improving 2.9%, 5.8%, and 2.7% over PPAC. PPD also outperforms the post-hoc baseline DAP (Chen et al., 2024), as PPD estimates interaction-level popularity for debiasing by considering both global and personalized preferences, rather than clustering similar nodes and relying on simple node degree information. Moreover, PPD consistently outperforms aggregation-weighting methods such as APDA (Zhou et al., 2023) and NAVIP (Kim et al., 2022), supporting our claim that applying weights in every training epoch can distort model learning and does not effectively reduce popularity bias. These results demonstrate several key insights. First, the substantial improvements on KuaiRec emphasize that PPD is particularly effective in settings where popularity can be most prominent, confirming its robustness to extreme skew in interaction frequency. Second, the consistent improvements on Coat indicate that PPD also generalizes well to datasets with less pronounced popularity bias. Finally, even on Yahoo! R3, where extremely sparse interactions can make debiasing particularly difficult due to fewer preference signals in the data, PPD yields consistent improvements, suggesting its ability to correct for bias in sparser datasets. Together, these findings show that PPD consistently outperforms the baselines across datasets with varying levels of skewness and sparsity.
In addition to performance, PPD demonstrates computational advantages over existing methods because it does not require retraining the model. As shown in Table 2, PPD is consistently faster than all methods across all datasets. It achieves 2.32×, 28.64×, and 3.44× speedups on KuaiRec, Coat, and Yahoo! R3, respectively, compared to the best-performing baselines. These results demonstrate that PPD not only improves recommendation performance by reducing popularity bias but also lowers computational cost. Moreover, PPD requires only the embeddings from the pre-trained model, avoiding the need to load and process the training data or train the GNN model on the GPU. As a result, it requires less GPU memory. PPD requires approximately 401MB, 228MB, and 352MB of GPU memory on KuaiRec, Coat, and Yahoo! R3, respectively, which is approximately half of that used by the best-performing baselines.
@@
Popular and niche item performance improvement (RQ2). RQ2 evaluates performance on popular (head) and niche (tail) items, examining whether improvements on one group come at the expense of the other. Popularity bias correction can create such a trade-off, where enhancing tail-item performance may degrade head-item accuracy (Zhou et al., 2023). To investigate this, following (Zhou et al., 2023; Chen et al., 2024), we divide items into the top 20% most frequently interacted ones (popular) and the remaining 80% (niche) based on the training data, and evaluate performance on these subsets. This allows assessment of whether PPD balances recommendation quality for both head and tail items. We choose the 20%-80% split, as it is commonly used in recommender systems to analyze popularity (Zhou et al., 2023; Chen et al., 2024), reflecting power-law popularity distributions. As shown in Table 3, PPD consistently achieves the best performance across both groups. On KuaiRec, PPD improves head performance by +186.7% in Recall@20 and +492.3% in NDCG@20 compared to the best baseline. For niche items, where all methods struggle due to very skewed data in tails, PPD still shows the best performance, achieving relative improvements of +233.3% in Recall@20 and +400% in NDCG@20 compared to best baseline. The overall low scores on KuaiRec tails stem from the extremely imbalanced distribution, where most niche items may have only a few interactions, making them inherently difficult to recommend. On Coat, PPD also improves both head and niche groups simultaneously. For head items, it yields +2.7% Recall@20 and +0.94% NDCG@20 over the best baseline. For niche items, it further improves by +1.24% Recall@20 and +1.09% NDCG@20, showing that gains on tail recommendations do not compromise head performance. On Yahoo! R3, PPD improves head recommendations with +7.1% Recall@20 and +12.0% NDCG@20 compared to best baseline, while also improving tail performance with +2.33% Recall@20, though NDCG shows a slight drop (–1.5%) compared to best baseline. Across all datasets, PPD achieves the highest HR@20 compared to the baselines, though we do not report the results due to space constraints. In summary, PPD enhances recommendations for tail items, demonstrating its ability to promote niche items while maintaining strong performance for head items. PPD explicitly removes popularity-driven components from embeddings while retaining preference-driven signals through interaction-level popularity scores. This ensures that head items are recommended when they genuinely reflect user interests, and tail items are recommended when aligned with user preferences. Therefore, the typical head–tail trade-off is avoided: head accuracy is preserved, and tail performance is improved.
@@
Effect of GNN layer depth on model performance (RQ3). We investigate how varying the number of GNN layers influences the performance of the PPD method. Figure 2 demonstrates the effect of increasing the number of layers. On the KuaiRec dataset, where popularity bias is most severe, LightGCN reaches peak performance with two layers but suffers degradation as more layers are added, indicating that deeper propagation amplifies popularity bias and over-smoothing. In contrast, PPD continues to benefit from additional layers, suggesting that its debiasing mechanism can mitigate bias amplification while leveraging multi-hop information. On the Coat dataset, where popularity bias is less pronounced, both models achieve their best performance with two layers, and deeper propagation beyond this point may introduce noise rather than useful signals, reducing accuracy. For Yahoo! R3, PPD achieves optimal results with two layers, while LightGCN slightly peaks at three layers, reflecting that in moderately biased and sparse settings, LightGCN can still gain from slightly deeper propagation. Importantly, across all datasets and all depths, PPD consistently outperforms LightGCN, demonstrating robustness and resilience to over-smoothing and bias mitigation. For NDCG@20, PPD also outperforms across different layer depths, exhibiting trends similar to Recall@20. These findings highlight that the optimal number of layers depends on the dataset, and PPD maintains stronger performance by mitigating popularity bias across different depths.
RQ4-RQ7. Beyond all these main results from RQ1-RQ3, we also evaluate four additional analysis in RQ4-RQ7. We present detailed results and analysis for RQ4–RQ7 in Appendix B.3. In RQ4, we apply PPD to another stronger GNN-based CF backbone, Self-supervised Graph Learning (SGL), to evaluate the generalizability of our debiasing strategy. PPD consistently outperforms SGL across all three datasets (see Table 4), with particularly large gains on the popularity-dominated KuaiRec dataset, demonstrating its effectiveness on different GNN-based CF backbones. In RQ5, we examine the role of personalized preference ($r_{ui}$) when estimating the popularity score ($b_{ui}$). The comparison between PPD and its variant without personalized preference (PPD w/o-PP) shows that removing this component when estimating the popularity score reduces performance across datasets (see Table 5), as relying solely on global preference risks discarding genuinely relevant popular items. This confirms that personalized preference is crucial for isolating the residual effect of popularity. In RQ6, we study the effect of the preference centroid coefficient ($\phi$) hyperparameter, which balances the contribution of preference signals in popularity direction. The results reveal that the optimal $\phi$ is dataset-dependent. Moderate values (e.g., $\phi=0.5$) work best when popularity is high and preference signals are weak, as in KuaiRec. Higher values yield better performance in less popularity-skewed datasets where preferences are more reliable, as in Coat and Yahoo! R3 (see Figure 3). These findings highlight the importance of tuning the preference centroid and adapting $\phi$ to each dataset for popularity debiasing. In RQ7, we examine the effect of the popularity penalty coefficient ($\beta$) on PPD during personalized preference estimation. On KuaiRec, Recall@20 improves as $\beta$ increases up to 0.2 but declines at 0.3, while Coat achieves its best performance at $\beta=0.1$ and Yahoo! R3 at $\beta=0.3$ (see Figure 4). The results also show that without a penalty ($\beta=0.0$), the model fails to reach optimal performance, highlighting the need for a popularity penalty in estimating personalized preferences.